OpenStudy (stuck-help):

Find the images vertices for a dilation with center(0,0) and a scale factor of 4.

OpenStudy (stuck-help):

@malcolmmcswain i dont know how to find vertices can you help

OpenStudy (malcolmmcswain):

This is a tricky one.

OpenStudy (malcolmmcswain):

I know an easy, but complicated way to do this.

OpenStudy (malcolmmcswain):

Wanna try it?

OpenStudy (stuck-help):

ok lets see if i can follow that way

OpenStudy (malcolmmcswain):

Never mind. I have a waaayyyyy easier way. But before I show you it, I need to explain how it works: Let's start off with this figure here: https://www.desmos.com/calculator/lttgxf6exs

OpenStudy (malcolmmcswain):

Oops sorry wrong link

OpenStudy (malcolmmcswain):

https://www.desmos.com/calculator/ue9blhj06i

OpenStudy (malcolmmcswain):

There's your figure.

OpenStudy (malcolmmcswain):

Now, let's take the point (4,-3)

OpenStudy (stuck-help):

ok

OpenStudy (malcolmmcswain):

And expand it outwards along a line, like this: https://www.desmos.com/calculator/kvjqstpl7y If the take the distance between the point and the origin (0,0) like this: 

OpenStudy (malcolmmcswain):

Are you following?

OpenStudy (stuck-help):

i cant read the drawing

OpenStudy (malcolmmcswain):

It says if the distance between the point and the origin is n, then the translated points distance from the origin will be 4 * n

OpenStudy (malcolmmcswain):

(if we are using a scale factor of 4)

OpenStudy (stuck-help):

ok i got that

OpenStudy (malcolmmcswain):

To find n we need x and y.

OpenStudy (stuck-help):

what is x and y

OpenStudy (malcolmmcswain):

So, if we let theta be the scale factor. The translated point will be: \[((x \times \theta),(y \times \theta)\]

OpenStudy (malcolmmcswain):

So, for a scale factor of 4, we can find each translated vertex with (x*4),(y*4)

OpenStudy (stuck-help):

im confused where do you get x and y

OpenStudy (malcolmmcswain):

x and y are what you use to get coordinates...

OpenStudy (stuck-help):

ok but what are referring to as x and y

OpenStudy (malcolmmcswain):

Ok. Let me help you visualize. Instead of theta, let's use s to represent the scale factor.

OpenStudy (stuck-help):

ok

OpenStudy (malcolmmcswain):

https://www.desmos.com/calculator/vjvbnc5dnv The green points represent the preimage. The blue points represent the dilated image. S represents the scale factor. When we change s, we change the dilation. Play around with it a bit until you understand.